The electrons in an electron microscope have speed of 4.7 Mm/s . Part A Find their de Broglie wavelength. Express your answer to two significant figures and include the appropriate units.
Answer #1

given

electron speed(v) = 4.7Mm/s

de broglie wavelength = sqrt( 2m KE )

m= mass of electron = 9.109*10^-31kg

KE = kinetic energy = 1/2*m*v^2

KE = 0.5*(9.109*10^-31)*(4.7^2)

KE = 1.006*10^-29 kg.Mm/s

de broglie wavelength = sqrt( 2m KE )

= sqrt(2*(9.109*10^-31kg)*(1.006*10^-29kg.Mm/s))

= 4.28*10^-30